[figure: 9.7 cm]

INTERLAYER COUPLING IN MAGNETIC SUPERLATTICES

WITH ELECTRON DENSITY INHOMOGENEITIES

W. Gruhn^{a} and Z. Ba̧k^{a} aInstitute of Physics, Pedagogical University, Al. Armii Krajowej 13/15, 42-200 Czȩstochowa, Poland

We discuss the influence of spatial inhomogeneities of the free electron density on the magnetic interaction between magnetic layers of the superlattice, mediated across nonmagnetic, metallic specer. Using the modified total energy approach, we prove that the TM or RE superlettices the additional scattering of free electrons on magnetic ion multipole moment increases the ferroquadrupolar biquadratic coupling between magnetic layers. We show that this novel mechanism, should manifest itself in the anisotropy of the magnetoresistivity.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : W. Gruhn

Address for correspondence : Institute of Physics, Pedagogical University, Al. Armii Krajowej 13/15, 42-200 Czȩstochowa, Poland

Email address : w.gruhn@wsp.czest.pl